Q2. If the consumption function of an economy is given by C=500+0.75Yd and Investment
function is I= 300, then:
a) Derive equation for AD.
b) Determine equilibrium value of output.
1
Expert's answer
2020-05-08T14:18:59-0400

If C = 500 + 0.75Yd and I = 300, then:

a) AD = Y = C + I = 500 + 0.75(Y - T) + 300 = 800 + 0.75(Y - T).

b) The equilibrium value of output is:

Y = 800 + 0.75(Y - T),

0.25Y = 800 - 0.75T,

Y = 3,200 - 3T, where T is for taxes.